# Mend the gap: A smart repair algorithm for noisy polygonal tilings

### v1

@@ 1.2. Problem statement
    For each $k=1,\ldots,N$, the area of $\widetilde{P}_{k}\cap P^{*}_{k}$ is as large as possble.
@@ 1.3. Outline of the smart_repair algorithm
    Nesting into larger units: In some applications, the polygons in $T$ are intended to nest cleanly into some larger units; e.g., in many states, voting precincts should nest cleanly into counties. The user may optionally specify a clean tiling $T^{\prime}$ of a region $R^{\prime}$ that closely approximates $R$—e.g., a map of county boundaries within a state—and then performs the repair process so that the repaired polygons in $\widetilde{T}$ nest cleanly into the polygons in $T^{\prime}$.
@@
While the development of the smart_repair algorithm was motivated by the author’s work in redistricting, its potential applications to versions of Problem 1.1 extend far beyond this context. Geographic Information Systems (GIS) are used for an enormous variety of applications, and geospatial data is often rife with gaps, overlaps, and other problems due to rounding errors and other inaccuracies in polygon boundaries. This issue seems to have received relatively little attention from the computational geometry community to date; we hope that this paper will spark further interest in geometric approaches to improving the quality of geospatial data.
@@ 2. Polygon geometry
  • A polygonal path $\overline{v_{1}v_{2}\cdots v_{k}}$ is a sequence of points $v_{1},v_{2},\ldots,v_{k}$ in the plane, called the vertices of the path, for which every pair of adjacent points $v_{i}$, $v_{i+1}$ ($11\leq i\leq k-$ 1) represents the line segment joining $v_{i}$ to $v_{i+1}$, and no two non-consecutive segments intersect.
@@
  • The exterior angle of a simple polygon $P$ at a vertex $v_{k}$ is the angle $\varphi=\pi-\theta$, where $\theta$ is the interior angle of $P$ at $v_{k}$. The exterior angle takes values in the range $-\pi<\varphi<\pi$.
@@
Note that the containment condition on the line segment $\overline{p^{\prime}p^{\prime\prime}}$ in Definition 2.5 is an open condition on the points $p^{\prime}\in B^{\prime}$, $p^{\prime\prime}\in B^{\prime\prime}$, so strong mutual visibility of a pair of paths $(B^{\prime},B^{\prime\prime})$ in $P$ implies that there exist nonempty open subsets $U^{\prime}\subset B^{\prime}$ and $U^{\prime\prime}\subset B^{\prime\prime}$ such that for every pair of points $(q^{\prime},q^{\prime\prime})$ with $q^{\prime}\in U^{\prime}$ and $q^{\prime\prime}\in U^{\prime\prime}$, $q^{\prime}$ and $q^{\prime\prime}$ are strongly visible to each other in $P$.
@@
Every simple polygon contains at least 3 convex vertices; this is an immediate consequence of the fact that every simple polygon has exterior angles that sum to $2\pi$, and each exterior angle is strictly less than $\pi$ and greater than 00 if and only if the corresponding vertex is convex. Moreover, since each of the paths $B_{1},\ldots,B_{m}$ is outward convex, the convex vertices of $P$ must be a subset of the intersection points $\{v_{1},\ldots,v_{m}\}$ between these paths.
Let $v_{k}$ be a convex vertex of $P$ with interior angle $\theta$. For $0<\alpha<\theta$, let $\ell_{\alpha}$ be the line segment in $P$ that starts at $v_{k}$, forms an interior angle of $\alpha$ with the edge of $\partial P$ emanating from $v_{k}$, and ends at the first point $q_{\alpha}\in\partial P$ where it intersects $\partial P$ at a positive distance from $v_{k}$. By construction, $\ell_{\alpha}$ is not contained in either $B_{k}$ or $B_{k+1}$, and $\ell_{\alpha}$ is the shortest path from $v_{k}$ to $q_{\alpha}$; therefore, Proposition 2.3 implies that $q_{\alpha}\notin B_{k}\cup B_{k+1}$. It follows that $q_{\alpha}\in B_{j_{\alpha}}$ for some $j_{\alpha}\neq k,k+1$. As $\alpha$ varies continuously between 00 and $\theta$, there must be some nonempty, open subinterval $(a,b)\subset(0,\theta)$ on which $j_{\alpha}$ is equal to a constant value $j_{0}$, and for which the points $\{q_{\alpha}\mid\alpha\in(a,b)\}$ form an open subset of $B_{j_{0}}$.
@@ 3.3.2. Gaps with 2 sub-boundaries
For gaps with 2 sub-boundaries, we accomplish this by constructing the shortest path within the gap between the endpoints of the sub-boundaries, dividing the gap along this path, and assigning the resulting regions to their adjacent units. Since the vertices in the shortest path are guaranteed to be vertices of the gap boundary (cf.  Lemma 2.2), this process does not introduce any new vertices or points of intersection between unit boundaries.666 Because the vast majority of gaps encountered in practice are simply connected and shortest paths within non-simply connected polygons are not guaranteed to be unique, the Maup implementation of smart_repair is restricted to simply connected gaps. Any non-simply connected gaps (e.g., large bodies of water containing islands) are left unfilled and the user is notified of their presence. (In practice, non-simply connected gaps are often large enough to remain unfilled by default in any case.) An example of this process is shown in Figure 15.
@@ 3.3.3. Gaps with 3 sub-boundaries
A straightforward application of the triangle inequality shows that the shortest paths between endpoints of different sub-boundaries cannot cross each other—although they may intersect at a vertex or along a polygonal path—and therefore the regions created in this way are guaranteed to intersect only along their boundaries (if at all), and only along entire boundary line segments and/or at isolated vertices. Therefore, this process does not create any overlaps or introduce any new vertices or points of intersection between unit polygon boundaries. We will refer to this process as convexification of the gap sub-boundaries.
Unlike in the case of gaps with 2 sub-boundaries, when there are 3 sub-boundaries convexification generally does not completely fill the gap. It does, however, leave a smaller gap with 3 sub-boundaries and a simpler geometric structure: Proposition 2.3 implies that after a gap has been simplified by convexifying its sub-boundaries, the new sub-boundaries of the simplified gap are all outward convex.
For simplified gaps with 3 sub-boundaries, we proceed as follows:
@@
    Construct the angle bisectors of the vertices where the endpoints of the sub-boundaries meet. The outward convexity of each sub-boundary implies that the angle bisector of the vertex where two sub-boundaries meet will necessarily intersect the gap boundary only at this vertex and at some point of the third sub-boundary. Consequently, each pair of angle bisectors will intersect at some interior point of the gap. The three interior points of intersection between the angle bisectors will either be coincident or form a triangle whose edges are subsegments of the angle bisectors.
@@
    If the intersection points are coincident, then the segments of the angle bisectors from their initial points to their common point of intersection suffice to partition the gap into regions that can each be assigned to their adjacent unit, similarly to the procedure for a simple triangle.
    If the intersection points are not coincident, then the segments of the angle bisectors from their initial points to the more distant points of intersection with the other angle bisectors partition the gap into 3 regions that are each adjacent to one of the units and the interior triangle formed by the intersection points. We assign each of the first 3 regions to its adjacent unit, leaving only the interior triangle unfilled. Finally, we use the procedure described above to fill the interior triangle, thereby completely filling the gap.
An example of this process applied to a gap with 3 sub-boundaries is shown in Figure 17.
Figure 17. (a) A gap with 3 sub-boundaries; (b) Shortest paths between sub-boundary endpoints within gap; (c) Simplified unit polygons and simplified gap; (d) Angle bisectors of vertices of simplified gap; (e) Partially repaired unit polygons and simple triangle gap; (f) Completely repaired unit polygons; (g) Completely filled gap
@@ 3.3.4. Gaps with 4 or more sub-boundaries
    • If $\alpha_{1}$ and $\alpha_{2}$ intersect, go back to the previous step and select the non-adjacent pair of sub-boundaries with the next-shortest distance between them.
    • If $\alpha_{1}$ and $\alpha_{2}$ are disjoint, then go on to the next step in order to create an adjacency between the unit polygons adjacent to $B^{\prime}$ and $B^{\prime\prime}$.
@@
    Construct the shortest path $\beta_{1}$ within the gap between the initial points of $B^{\prime}$ and $B^{\prime\prime}$ and the shortest path $\beta_{2}$ within the gap between the terminal points of $B^{\prime}$ and $B^{\prime\prime}$. By Theorem 2.11, the union of $B^{\prime}$, $B^{\prime\prime}$, $\beta_{1}$, and $\beta_{2}$ bounds either one or two simple polygons, and each of these polygons intersects exactly one of $B^{\prime}$, $B^{\prime\prime}$ in a path of positive length.
@@
    Assign each polygon created in the previous step to the unit polygon with which it shares a boundary of positive length. According to Theorem 2.11, the resulting unit polygons will now intersect at a single common vertex. Furthermore, this will leave either one or two smaller gaps remaining, each of which contains strictly fewer sub-boundaries than the original gap.
@@ 4.2. Small rook-to-queen adjacency conversion
In rare cases, it may happen that two or more of the disks constructed in Step (1) above intersect nontrivially. In this case, we take the convex hull of the union of these disks and apply the remainder of the procedure to this polygon, so that all of the small rook adjacencies within this polygon are converted to a single queen adjacency. For example, when we applied this procedure to the Colorado 2020 voting precinct map with a length threshold of 100 feet, we found 193 adjacencies below the threshold, including two pairs for which the disks overlapped. One of these “double disks” and its intersection with 5 precincts, together with the result of the conversion procedure, is shown in Figure 23.
@@ 5. Runtime complexity
In this section we perform a rough estimate for the runtime complexity for the smart_repair algorithm. This estimate is not intended to be optimal, and it is not our primary concern, as the development of the algorithm was motivated primarily by a desire for accuracy rather than efficiency. We have found that in practice, the Maup implementation can repair most state-level precinct maps in a few hours on a 2019 Intel-based MacBook Pro, which is more than adequate for our needs.
@@
  • The fully noded union of the polygon boundaries is constructed by computing intersection points between all edges in all polygons. This requires at most (${E\choose 2}$ operations (and in practice, generally much fewer than this [^12] For instance, the Bentley-Ottmann algorithm of [2] has run time complexity $O((E+k)\log E)$, where $k$ is the number of intersection points.), and the number $E^{\prime}$ of edges in the resulting fully noded union is bounded by $E^{\prime}\leq 4{E\choose 2}=2E(E-1)$.
@@
  • Remaining simplified gaps all have at least 3 sub-boundaries. For each gap with $s\geq 4$ sub-boundaries, it generally requires $(s-3)$ applications of the procedure outlined in §3.3.4 to reduce the gap to $(s-2)$ smaller gaps with 3 sub-boundaries each. Each application of this procedure requires computing the distance between each of the non-adjacent pairs of sub-boundaries in the gap, and then construction of 4 shortest paths between endpoints of the pair with the shortest distance, each with runtime complexity bounded by $O(E^{\prime}\log E^{\prime})$. Since the total number of sub-boundaries for all gaps is bounded by $E^{\prime}$, the total run time complexity for this step is at most $O((E^{\prime})^{3})$ to compute all distances between pairs of sub-boundaries for all $(s-3)$ iterations, and $O((E^{\prime})^{2}\log E^{\prime})$ for computing shortest paths, for an overall runtime complexity of $O((E^{\prime})^{3})$.
@@ 6. Conclusion
Finally, we present some statistics regarding gaps, overlaps, and the performance of the Maup implementation of smart_repair for a few representative state-level precinct maps. All computations were performed on a 2019 Intel-based MacBook Pro running Python 3.11 and using Maup 2.0.0.
  • The 2020 Colorado voting precinct map [^13] Compiled and provided by the staff of the Colorado Independent Legislative Redistricting Commission has 3,215 precincts, and the original map contained 909 overlaps and 1,475 gaps. [^14] Counts of gaps and overlaps were computed using the doctor function in the Maup package. The runtime for smart_repair function on this map was about 22 minutes.
  • The 2023 Wisconsin ward map [^15] Downloaded from the Wisconsin Legislative Technology Services Bureau at https://gis-ltsb.hub.arcgis.com/ has 7,013 wards, and the original map contained 4,148 overlaps and 10,881 gaps. The runtime for smart_repair function on this map was slightly over 7 hours.
  • The 2020 New York voting precinct map [^16] Compiled by the Voting and Elections Science Team (VEST) and downloaded from the Redistricting Data Hub at https://redistrictingdatahub.org has 15,376 precincts, and the original map contained 3,069 overlaps and 7,524 gaps. The runtime for smart_repair function on this map was about 23 hours.

### v2

@@ 1.2. Problem statement
    For each $k=1,\ldots,N$, the area of $\widetilde{P}_{k}\cap P^{*}_{k}$ is as large as possible.
@@ 1.3. Outline of the smart_repair algorithm
    Nesting into larger units: In some applications, the polygons in $T$ are intended to nest cleanly into some larger units; e.g., in many states, voting precincts should nest cleanly into counties. The user may optionally specify a clean tiling $T^{\prime}$ of a region $R^{\prime}$ that closely approximates RR—e.g., a map of county boundaries within a state—and then performs the repair process so that the repaired polygons in $\widetilde{T}$ nest cleanly into the polygons in $T^{\prime}$.
@@
While the development of the smart_repair algorithm was motivated by the author’s work in redistricting, its potential applications to versions of Problem 1.1 extend far beyond this context. Geographic Information Systems (GIS) are used for an enormous variety of applications, and geospatial data is often rife with gaps, overlaps, and other problems due to rounding errors and other inaccuracies in polygon boundaries. As just one example, in [6] the authors describe similar issues that arose in a project involving topological data analysis of COVID-19 data by ZIP code. As such, we hope that this algorithm and its Python implementation as part of the Maup package (see Footnote 3) will be of interest to the broader GIS community.
@@ 2. Polygon geometry
  • A polygonal path $\overline{v_{1}v_{2}\cdots v_{k}}$ is a sequence of points $v_{1},v_{2},\ldots,v_{k}$ in the plane, called the vertices of the path, for which every pair of adjacent points $v_{i}$, $v_{i+1}$ ($1\leq i\leq k-1$) represents the line segment joining $v_{i}$ to $v_{i+1}$, and no two non-consecutive segments intersect.
@@
  • The exterior angle of a simple polygon $P$ at a vertex $v_{k}$ is the angle $\varphi=\pi-\theta$, where $\theta$ is the interior angle of $P$ at $v_{k}$. The exterior angle takes values in the range $\pi<\varphi<\pi$.
@@
Note that the containment condition on the line segment $\overline{p^{\prime}p^{\prime\prime}}$ in Definition 2.5 is an open condition on the points $p^{\prime}\in B^{\prime}$, $p^{\prime\prime}\in B^{\prime\prime}$, so strong mutual visibility of a pair of paths $(B^{\prime},B^{\prime\prime})$ in $P$ implies that there exist nonempty open subsets $U^{\prime}\subset B^{\prime}$ and $U^{\prime\prime}\subset B^{\prime\prime}$ such that for every pair of points $(q^{\prime},q^{\prime\prime})$ with $q^{\prime}\in U^{\prime}$ and $^{\prime\prime}\in U^{\prime\prime}, qq^{\prime}$ and $q^{\prime\prime}$ are strongly visible to each other in $P$.
@@
Every simple polygon contains at least 3 convex vertices; this is an immediate consequence of the fact that every simple polygon has exterior angles that sum to $2\pi$, and each exterior angle is strictly less than $\pi$ and greater than 0 if and only if the corresponding vertex is convex. Moreover, since each of the paths $B_{1},\ldots,B_{m}$ is outward convex, the convex vertices of $P$ must be a subset of the intersection points $\{v_{1},\ldots,v_{m}\}$ between these paths.
Let $v_{k}$ be a convex vertex of $P$ with interior angle $\theta$. For $0<\alpha<\theta$, let $\ell_{\alpha}$ be the line segment in $P$ that starts at $v_{k}$, forms an interior angle of $\alpha$ with the edge of $\partial P$ emanating from $v_{k}$, and ends at the first point $q_{\alpha}\in\partial P$ where it intersects $\partial P$ at a positive distance from $v_{k}$. By construction, $\ell_{\alpha}$ is not contained in either $B_{k}$ or $B_{k+1}$, and $\ell_{\alpha}$ is the shortest path from $v_{k}$ to $q_{\alpha}$; therefore, Proposition 2.3 implies that $q_{\alpha}\notin B_{k}\cup B_{k+1}$. It follows that $q_{\alpha}\in B_{j_{\alpha}}$ for some $j_{\alpha}\neq k,k+1$. As $\alpha$ varies continuously between 0 and $\theta$, there must be some nonempty, open subinterval $(a,b)\subset(0,\theta)$ on which $j_{\alpha}$ is equal to a constant value $j_{0}$, and for which the points $\{q_{\alpha}\mid\alpha\in(a,b)\}$ form an open subset of $B_{j_{0}}$.
@@ 3.3.2. Gaps with 2 sub-boundaries
For gaps with 2 sub-boundaries, we accomplish this by constructing the shortest path within the gap between the endpoints of the sub-boundaries, dividing the gap along this path, and assigning the resulting regions to their adjacent units. Since the vertices in the shortest path are guaranteed to be vertices of the gap boundary (cf.  Lemma 2.2), this process does not introduce any new vertices or points of intersection between unit boundaries. [^6] Because the vast majority of gaps encountered in practice are simply connected and shortest paths within non-simply connected polygons are not guaranteed to be unique, the Maup implementation of smart_repair is restricted to simply connected gaps. Any non-simply connected gaps (e.g., large bodies of water containing islands) are left unfilled and the user is notified of their presence. (In practice, non-simply connected gaps are often large enough to remain unfilled by default in any case.) An example of this process is shown in Figure 15.
@@ 3.3.3. Gaps with 3 sub-boundaries
A straightforward application of the triangle inequality shows that the shortest paths between endpoints of different sub-boundaries cannot cross each other—although they may intersect at a vertex or along a polygonal path—and therefore the regions created in this way are guaranteed to intersect only along their boundaries (if at all), and only along entire boundary line segments and/or at isolated vertices. Therefore, this process does not create any overlaps or introduce any new vertices or points of intersection between unit polygon boundaries. We will refer to this process as convexification of the gap sub-boundaries; see Figure 17 for an example.
Figure 17. (a) A gap with 3 sub-boundaries; (b) Shortest paths within gap between sub-boundary endpoints; (c) Simplified unit polygons and simplified gap
Unlike in the case of gaps with 2 sub-boundaries, when there are 3 sub-boundaries convexification generally does not completely fill the gap. It does, however, leave a smaller gap with 3 sub-boundaries and a simpler geometric structure: Proposition 2.3 implies that after a gap has been simplified by convexifying its sub-boundaries, the new sub-boundaries of the simplified gap are all outward convex. This implies that the convex hull of the gap is a triangle whose vertices are the endpoints of the gap sub-boundaries. Let $q$ be the incenter of this triangle; then there are two possibilities:
@@
    If $q$ is contained in the interior of the gap, construct the shortest paths within the gap from $q$ to each of the endpoints of the sub-boundaries. These paths partition the gap into regions that can each be assigned to their adjacent unit, similarly to the procedure for a simple triangle. (See Figure 18 for an example.)
@@
    If $q$ is not contained in the interior of the gap, then it is contained in a polygon bounded by exactly one of the gap sub-boundaries and the boundary of the gap’s convex hull. Without loss of generality, call this sub-boundary $B_{1}$ and the other two sub-boundaries $B_{2}$ and $B_{3}$. Let $v_{0}$ be the gap vertex where $B_{2}$ and $B_{3}$ intersect, and let $v_{1}$ be the interior vertex of $B_{1}$ that is closest to $v_{0}$. (Note that this scenario implies that $B_{1}$ contains at least one interior vertex in addition to its endpoints.) Construct the shortest path within the gap between $v_{0}$ and $v_{1}$; this path partitions the gap into two regions, one adjacent to each of $B_{2}$ and $B_{3}$, which can then be assigned accordingly. (See Figure 19 for an example.)
Figure 18. (a) A gap with 3 convexified sub-boundaries; (b) Convex hull of the gap, with incenter contained in the interior of the gap; (c) Shortest paths within gap between incenter and sub-boundary endpoints; (d) Repaired unit polygons; (e) Filled gap
Figure 19. (a) A gap with 3 convexified sub-boundaries; (b) Convex hull of the gap, with incenter not contained in the interior of the gap; (c) Shortest path within gap between opposite vertex and nearest point on sub-boundary; (d) Repaired unit polygons; (e) Filled gap
@@ 3.3.4. Gaps with 4 or more sub-boundaries
    • If $\alpha_{1}$ and $\alpha_{2}$ intersect, then Theorem 2.9 implies that $B^{\prime}$ and $B^{\prime\prime}$ are not strongly mutually visible, so go back to the previous step and select the non-adjacent pair of sub-boundaries with the next-shortest distance between them.
    • If $\alpha_{1}$ and $\alpha_{2}$ are disjoint, then Theorem 2.9 implies that $B^{\prime}$ and $B^{\prime\prime}$ are strongly mutually visible, so go on to the next step in order to create an adjacency between the unit polygons adjacent to $B^{\prime}$ and $B^{\prime\prime}$.
@@
    Construct the shortest path $\beta_{1}$ within the gap between the initial points of $B^{\prime}$ and $B^{\prime\prime}$ and the shortest path $\beta_{2}$ within the gap between the terminal points of $B^{\prime}$ and $B^{\prime\prime}$. By Theorem 2.11, the union of $^{\prime}, B′′B^{\prime\prime}$, $\beta_{1}$, and $\beta_{2}$ bounds either one or two simple polygons, and each of these polygons intersects exactly one of $^{\prime}, B′′B^{\prime\prime}$ in a path of positive length.
@@
    Assign each polygon created in the previous step to the unit polygon adjacent to either $B^{\prime}$ or $B^{\prime\prime}$ with which it shares a boundary of positive length. According to Theorem 2.11, the resulting unit polygons will now intersect at a single common vertex. Furthermore, this will leave either one or two smaller gaps remaining, each of which contains strictly fewer sub-boundaries than the original gap.
@@ 4.2. Small rook-to-queen adjacency conversion
In rare cases, it may happen that two or more of the disks constructed in Step (1) above intersect nontrivially. In this case, we take the convex hull of the union of these disks and apply the remainder of the procedure to this polygon (with the polygon’s centroid replacing the center of the disk in the previous construction), so that all of the small rook adjacencies within this polygon are converted to a single queen adjacency. For example, when we applied this procedure to the Colorado 2020 voting precinct map with a length threshold of 100 feet, we found 188 adjacencies below the threshold, including two pairs for which the disks overlapped. One of these “double disks” and its intersection with 5 precincts, together with the result of the conversion procedure, is shown in Figure 25.
@@ 5. Runtime complexity
In this section we perform a rough estimate for the runtime complexity for the smart_repair algorithm. This estimate is not intended to be optimal, and it is not our primary concern, as the development of the algorithm was motivated primarily by a desire for accuracy rather than efficiency. We have found that in practice, the Maup implementation can repair most state-level precinct maps in a few hours on a standard laptop computer, which is more than adequate for our needs.
@@
  • The fully noded union of the polygon boundaries is constructed by computing intersection points between all edges in all polygons. This requires at most $E\choose 2$} operations (and in practice, generally much fewer than this [^12] For instance, the Bentley-Ottmann algorithm of [2] has run time complexity $O((E+k)\log E)$, where $k$ is the number of intersection points.), and the number $E^{\prime}$ of edges in the resulting fully noded union is bounded by $E^{\prime}\leq 4{E\choose 2}=2E(E-1)$.
@@
  • Remaining simplified gaps all have at least 3 sub-boundaries. For each gap with $s\geq 4$ sub-boundaries, it generally requires (s3)(s-3) applications of the procedure outlined in §3.3.4 to reduce the gap to (s2)(s-2) smaller gaps with 3 sub-boundaries each. Each application of this procedure requires computing the distance between each of the non-adjacent pairs of sub-boundaries in the gap, and then construction of 4 shortest paths between endpoints of the pair with the shortest distance, each with runtime complexity bounded by $O(E^{\prime}\log E^{\prime})$. Since the total number of sub-boundaries for all gaps is bounded by $E^{\prime}$, the total run time complexity for this step is at most $O((E^{\prime})^{3})$ to compute all distances between pairs of sub-boundaries for all (s3)(s-3) iterations, and $O((E^{\prime})^{2}\log E^{\prime})$ for computing shortest paths, for an overall runtime complexity of $O((E^{\prime})^{3})$.
@@
As for practical performance, here are some sample statistics regarding gaps, overlaps, and the performance of the Maup implementation of smart_repair for a few representative state-level precinct maps. All computations were performed on a 2023 MacBook Pro with M3 Max chip running Python 3.12 and using Maup 2.0.3.
  • The 2020 Colorado voting precinct map [^13] Compiled and provided by the staff of the Colorado Independent Legislative Redistricting Commission has 3,215 precincts, and the original map contained 909 overlaps and 1,475 gaps. [^14] Counts of gaps and overlaps were computed using the doctor function in the Maup package. The runtime for smart_repair function on this map was about 10 minutes.
  • The 2023 Wisconsin ward map [^15] Downloaded from the Wisconsin Legislative Technology Services Bureau at https://gis-ltsb.hub.arcgis.com/ has 7,013 wards, and the original map contained 4,109 overlaps and 10,821 gaps. The runtime for smart_repair function on this map was slightly under 4 hours.
  • The 2020 New York voting precinct map [^16] Compiled by the Voting and Elections Science Team (VEST) and downloaded from the Redistricting Data Hub at https://redistrictingdatahub.org has 15,376 precincts, and the original map contained 3,036 overlaps and 7,699 gaps. The runtime for smart_repair function on this map was slightly under 13 hours.
@@ 6. Conclusion
7. Data Availability Statement
The source code for the smart_repair function is available as part of the Maup package for Python at https://github.com/mggg/maup. The Colorado 2020 precinct shapefile is available from the author on reasonable request. The Wisconsin 2023 ward shapefile was obtained from https://gis-ltsb.hub.arcgis.com/ but has since been replaced by a newer version on that website; it is available from the author on reasonable request. The New York 2020 precinct shapefile is available either at https://redistrictingdatahub.org or https://dataverse.harvard.edu/dataverse/electionscience.
8. Funding